Two surveyors with two-way radios leave the same point at 9:00 A.M., one walking due south at 2 mi/hr and the other due west at 4 mi/hr . How long can they communicate with one another if each radio has a maximum range of 4 miles?
Answer by mananth(16949)

distance of west bound = 4*t = 4t miles t = time
distance of south bound = 2*t = 2t miles
..
Their positions after time t form a right triangle with the starting point.
8^2 = (4t)^2+(2t)^2
64= 16t^2+4t^2
64=20t^2
64/20 = t^2
16/5 = t^2
take square root
4/sqrt5 = t
t= 1.78 hours